OpenStudy (anonymous):

(cos^2/sin^2) + csc θ sin θ

OpenStudy (chaise):

What are we doing with this problem?

OpenStudy (jdoe0001):

so... what are we doing here? do we print it out and engrave it? use it as wallpaper picture?

OpenStudy (anonymous):

Use basic identities to simplify the expression

OpenStudy (chaise):

Well csc(x)=1/sin(x) and, tan(x)=sin(x)/cos(x) and, cot(x)=1/tan(x)=cos(x)/sin(x) Do these rules help, or do you need more help?

OpenStudy (anonymous):

I know csc(x)sin(x) = 1, so how do I simplify cos^2sin^2?

OpenStudy (loser66):